# atomic orbital API Application Programmer Interface

### 06 Mar 2024 v3

@@ 3.1 The generalized mean-field potential
$[\beta,M]_{+}=2\begin{pmatrix}M^{L_{\alpha}L_{\alpha}}&M^{L_{\alpha}L_{\beta}}&0&0\\ M^{L_{\beta}L_{\alpha}}&M^{L_{\beta}L_{\beta}}&0&0\\ 0&0&-M^{S_{\alpha}S_{\alpha}}&-M^{S_{\alpha}S_{\beta}}\\ 0&0&-M^{S_{\beta}S_{\alpha}}&-M^{S_{\beta}S_{\beta}}\end{pmatrix}.$ (25)
@@
The three terms in Eq. (21) enter at different orders in $\frac{1}{c}$. A possible performance gain, especially when working with larger molecular systems, would be to apply them with a threshold obtained by scaling the user-defined one with their respective denominators.
@@
For closed-shell, two-electron systems, we can explicitly exploit Kramers’ time-reversal symmetry (TRS) and remove the exchange terms:[46]
@@ 4 Results and Discussion
All MW results presented were obtained with our Python implementation in the ReMRChem project, whose source code is freely avaialble through GitHub. [^2] https://github.com/MRChemSoft/ReMRChem. Our implementation uses the MRCPP library [47] and its Python interface provided by the VAMPyR package [48, 49]. We have set up our calculations choosing the smeared nuclear point-like charge, which replaces the nuclear $1/r$ potential with a smoothed approximation[50]. This smoothing has no physical meaning and it must not be misunderstood with the physical models introduced by Visscher [51], which are common in relativistic methods. The value of the speed of light was chosen equal to 137.03599913900001 a.u. and for the calculation of the derivative we used the B-spline method introduced by Anderson et al.[52].
@@
The current implementation uses the VAMPyR package (Python code) and is a proof-of-concept implementation to validate the use of the squared Dirac operator within the multiwavelets framework. This implementation can only deal with at most two-electron systems since it relies, as already stated, on Kramers’ TRS [46, 53] to reduce memory footprint and computational load. This is the ideal starting point to consider inter-electronic interactions in the SCF setting for the validation of the developed theory. However, as it is known from the nonrelativistic MW implementation in the MRChem package (written in C++), increasing the number of electrons will not pose additional theoretical challenges, but only increase the required computational resources (memory in particular) as tested by our group and extensively discussed in Ref. 54.
@@
Our results show that at low precision the two implementations are equivalent, when precision is increased, the $\hat{\mathfrak{D}}^{2}$ implementation leads to slightly lower errors (up to one order of magnitude) than the $\hat{\mathfrak{D}}$ implementation.
@@
$\hat{\mathfrak{D}}(\epsilon=10^{-4})$0.500007467791874 $|1.62236|\cdot 10^{-06}$1.102640862147382 $|6.51808|\cdot 10^{-07}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-4})$0.500007785136819 $|2.25705|\cdot 10^{-06}$1.102640727757642 $|7.73929|\cdot 10^{-07}$
$\hat{\mathfrak{D}}(\epsilon=10^{-6})$0.500006653182936 $|6.83203|\cdot 10^{-09}$1.102641514193238 $|6.06142|\cdot 10^{-08}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-6})$0.500006655675723 $|1.84653|\cdot 10^{-09}$1.102641581857073 $|7.47750|\cdot 10^{-10}$
$\hat{\mathfrak{D}}(\epsilon=10^{-8})$0.500006656217010 $|7.63966|\cdot 10^{-10}$1.102641581459466 $|3.77258|\cdot 10^{-10}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-8})$0.500006656559709 $|7.85770|\cdot 10^{-11}$1.102641580981426 $|4.63886|\cdot 10^{-11}$
Exact result -0.500006656598998 -1.102641581032577
@@
22footnotetext: Legendre polynomial order $k=10$ and tolerance $\epsilon=10^{-8}$, Legendre polynomial order $k=8$ and tolerance $\epsilon=10^{-6}$, Legendre polynomial order $k=6$ and tolerance $\epsilon=10^{-4}$.
@@
We report the comparison of our results with radial numerical results previously obtained using GRASP Ref. 35 in Tab. 2. Also in this case we have performed calculations by increasing requested precision both for the $\hat{\mathfrak{D}}$ and $\hat{\mathfrak{D}}^{2}$ formulations. We observe that both implementations are able to achieve the requested precision. However, there is no noticeable difference between them.
@@
Operator orbital energy Abs Rel. Error Total energy Abs Rel. Error
VAMPyR / (Ha) VAMPyR / (Ha)
$\hat{\mathfrak{D}}(\epsilon=10^{-4})$0.918040748 $|5.45383|\cdot 10^{-05}$2.861854359 $|1.43330|\cdot 10^{-05}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-4})$0.918032833 $|4.59170|\cdot 10^{-05}$2.861828187 $|5.18788|\cdot 10^{-06}$
$\hat{\mathfrak{D}}(\epsilon=10^{-6})$0.918008207 $|1.90908|\cdot 10^{-05}$2.861833848 $|7.16586|\cdot 10^{-06}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-6})$0.918008259 $|1.91475|\cdot 10^{-05}$2.861833976 $|7.21072|\cdot 10^{-06}$
$\hat{\mathfrak{D}}(\epsilon=10^{-8})$0.917990690 $|8.78784|\cdot 10^{-09}$2.861813340 $|1.93716|\cdot 10^{-10}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-8})$0.917990691 $|9.69563|\cdot 10^{-09}$2.861813342 $|5.19185|\cdot 10^{-10}$
GRASP Reference -0.917990682 -2.861813340
[…]
22footnotetext: Legendre polynomial order $k=10$ and tolerance $\epsilon=10^{-8}$, Legendre polynomial order $k=8$ and tolerance $\epsilon=10^{-6}$, Legendre polynomial order $k=6$ and tolerance $\epsilon=10^{-4}$.
[…]
Operator orbital energy Abs Rel. Error Total energy Abs Rel. Error
VAMPyR / (Ha) VAMPyR / (Ha)
$\hat{\mathfrak{D}}(\epsilon=10^{-4})$44.094786197 $|2.794189|\cdot 10^{-03}$94.108543063 $|1.337941|\cdot 10^{-03}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-4})$43.978792210 $|1.562788|\cdot 10^{-04}$93.990113164 $|7.781809|\cdot 10^{-05}$
$\hat{\mathfrak{D}}(\epsilon=10^{-6})$43.976907059 $|1.134071|\cdot 10^{-04}$93.987884382 $|5.410330|\cdot 10^{-05}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-6})$43.971930116 $|2.224998|\cdot 10^{-07}$93.982819929 $|2.162733|\cdot 10^{-07}$
$\hat{\mathfrak{D}}(\epsilon=10^{-8})$43.971921492 $|2.638859|\cdot 10^{-08}$93.982804435 $|5.142152|\cdot 10^{-08}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-8})$43.971921513 $|2.686178|\cdot 10^{-08}$93.982804554 $|5.268882|\cdot 10^{-08}$
GRASP Reference -43.971920332 -93.982799603
[…]
22footnotetext: Legendre polynomial order $k=10$ and tolerance $\epsilon=10^{-8}$, Legendre polynomial order $k=8$ and tolerance $\epsilon=10^{-6}$, Legendre polynomial order $k=6$ and tolerance $\epsilon=10^{-4}$.
[…]
5 Conclusions
The implementation of the squared Dirac operator $\hat{\mathfrak{D}}^{2}$ within a MW framework allows the use of the same integral equation formalism which has been widely tested in the context of non-relativistic theory. The original papers reporting the use of such operators [4, 31], indicated the importance of making use of a large (by practical means complete) basis set to achieve valid results with this method. The main reason is to make sure that the matrix representation of the squared operator is equivalent to the matrix product in Eq. (23). Such a requirement was challenging to satisfy in an atom-centered basis with the limited computational resources available in the early 1980s, and to some extent it is not trivial to achieve for Gaussian Type Orbitals. On the other hand, MWs have by and large demonstrated that they can be regarded as a complete basis set.
In the present work, we have shown that the use of the $\hat{\mathfrak{D}}^{2}$ operator with MWs is possible and it provides results with an error with respect to radial numerical approach (i.e., GRASP) of the same order of magnitude as using the $\hat{\mathfrak{D}}$ operator.

### 27 Feb 2026 v4

@@ 3.1 The generalized mean-field potential
$\beta,M]_{+}=2\begin{pmatrix}M^{L_{\alpha}L_{\alpha}}&M^{L_{\alpha}L_{\beta}}&0&0\\ M^{L_{\beta}L_{\alpha}}&M^{L_{\beta}L_{\beta}}&0&0\\ 0&0&-M^{S_{\alpha}S_{\alpha}}&-M^{S_{\alpha}S_{\beta}}\\ 0&0&-M^{S_{\beta}S_{\alpha}}&-M^{S_{\beta}S_{\beta}}\end{pmatrix}$. (25)
@@
For closed-shell, two-electron systems, we can explicitly exploit Kramers’ time-reversal symmetry (TRS) and remove the exchange terms:[Kramers1930-hq]
@@
3.2 Cross expectation values
We can use two different approaches to converge our result toward the ground state: the Dirac-Fock integral equation (18c), or the Squared Dirac Operator integral equation (22). Both algorithms will effectively evolve the initial guess spinor function towards the ground state energy solutions, by propagating the component functions in the Hilbert space until the total update will be lower than the set threshold [Beylkin2021-sk]. Once convergence is reached, we can evaluate the final ground state energy either by computing the expectation value of $\hat{\mathfrak{D}}$ or $\hat{\mathfrak{D}}^{2}$.
To better discuss the methods and the results, we will denote by $\varepsilon(\hat{\mathcal{O}}_{E},\Phi_{\hat{\mathcal{O}}_{SCF}})$ the orbital energy obtained by taking the expectation value of $\hat{\mathcal{O}}_{E}$ for the wavefunction $\Phi_{\hat{\mathcal{O}}_{SCF}}$, obtained by performing the SCF using the $\hat{\mathcal{O}}_{SCF}$ operator. We will thus have four different possibilities:
$\begin{pmatrix}\varepsilon(\hat{{\mathfrak{D}}},\Phi_{\hat{{\mathfrak{D}}}})&\varepsilon(\hat{{\mathfrak{D}}}^{2},\Phi_{\hat{{\mathfrak{D}}}})\\ \varepsilon(\hat{{\mathfrak{D}}},\Phi_{\hat{{\mathfrak{D}}}^{2}})&\varepsilon(\hat{{\mathfrak{D}}}^{2},\Phi_{\hat{{\mathfrak{D}}}^{2}})\end{pmatrix}$ (31)
The diagonal elements represent results where both the wave function optimization and the energy expectation value are obtained with the same method. The off-diagonal terms represent results where the wave function optimization and the energy are calculated using different methods. This adds a further dimension of variables to investigate that could lead to more precise result by using less resources.
3.3 Amplification of relativistic effects for lighter atoms
It is sometimes convenient to be able to run calculation with an artificial speed of light. For higher values one expects to get non-relativistic results, whereas for lower values one can amplify relativistic effects. The latter is particularly useful here to understand if the observed precision (or lack thereof) should be attributed to the relativistic effect or to the nuclear charge. For example, assume we want to achieve the same relativistic correction for Ar as one would observe for Hg. We can impose that the new speed of light, when inserted in the analytical formula for the 1-electron atom in the case of a $1s$ electron for Argon, would yield the same relative energy shift as in Mercury with the physically correct $c$. Given the analytical solution for the ground state of a single electron system for the Dirac equation in atomic units:
$E^{R}(c)=c^{2}\sqrt{1-\frac{Z^{2}}{c^{2}}}-c^{2},$ (32)
, we want to obtain the following:
$\frac{E^{NR}_{{}\mathrm{Ar}}-E^{R}_{{}\mathrm{Ar}}(c^{\prime})}{E^{NR}_{{}\mathrm{Ar}}}=\frac{E^{NR}_{{}\mathrm{Hg}}-E^{R}_{{}\mathrm{Hg}}(c)}{E^{NR}_{{}\mathrm{Hg}}}$ (33)
This is verified if $c^{\prime}=\frac{Z_{{}\mathrm{Ar}}}{Z_{{}\mathrm{Hg}}}\;c$.
@@ 4 Results and Discussion
The goal of the numerical tests presented here is to asses the precision of $\hat{\mathfrak{D}}^{2}$ with respect to $\hat{\mathfrak{D}}$ in the MW framework. As a reference we have taken the results obtained with the GRASP[grasp_ref] code. The tests will investigate the precision obtained for both methods, employed both in the SCF optimization and to compute the orbital energy. This yields a total of four combination as outlined in Eq. (31). We also assessed the effects of the choice of derivative operator ( B-Spline (BS[ANDERSON2019100033] vs. Alpert, Beylkin, Ginez and Vozovoi (ABGV[Alpert_MRA] realizations) and grid choice (nuclear charge on a dyadic point or offset from it).
[…]
Except when specified otherwise, the speed of light was set to $c=137.0359895\;\textit{a.u.}$. For the two smallest nuclei (H and He) we made use of a simple point-like charge model. For the MW calculations the usual numerical, precision-dependent smoothing described in [Yanai2004] was employed, to avoid an excessive refinement of the grid when trying to represent the potential of the nucleus. For heavier nuclei (Ne, Ar, Hg) we employed a Fermi-Dirac charge distribution [Visscher_1997] as this is the model available in the GRASP[grasp_ref] software as we have done in a previous study [tantardini2023future]. In Table 1 we report the explicit parameters employed in the two codes. The difference is due to the conventions adopted. In ReMRChem[ReMRChem], the skin thickness is defined as described in [Visscher_1997], with a constant a=2.3a=2.3fm (then converted in a.u.), while the half charge radius varies from atom to atom. GRASP[grasp_ref] uses the same physical model but it employs a thickness parameter rescaled by the factor $4\ln{(3)}$, which appears explicitly in the original article [Visscher_1997]. This is why the $a$ parameter is different in the two inputs. We emphasize this technical point to ease reproducibility and because it required quite some effort to track down this subtlety when comparing our results to GRASP[grasp_ref].
Table 1: All the values are given and used in atomic units (Bohr radii).
Grasp VAMPyR
nucleus $a$ $c$ $a$ $c$
Ne $9.89059\,10^{-6}$ $5.589069\,10^{-5}$ $4.34637\,10^{-5}$ $5.589069\,10^{-5}$
Ar $9.89059\,10^{-6}$ $6.883946\,10^{-5}$ $4.34637\,10^{-5}$ $6.883946\,10^{-5}$
Hg $9.89059\,10^{-6}$ $1.2433249\,10^{-4}$ $4.34637\,10^{-5}$ $1.2433249\,10^{-4}$
@@
All MW results presented were obtained with our Python implementation in the ReMRChem[ReMRChem] project, whose source code is freely available through GitHub. Our implementation uses the mrcpp[bast_2023_7967323] library and its Python interface provided by the VAMPyR[battistella_2023_10290360] package [bjorgve2024]. This implementation can only deal with at most two-electron systems since it relies, as already stated, on Kramers’ TRS [Kramers1930-hq, Wigner_1932] to reduce memory footprint and computational cost and can be considered a proof-of-concept implementation to validate the use of the squared Dirac operator within the MW framework. This is the ideal starting point to consider inter-electronic interactions in the SCF setting for the validation of the developed theory. However, as it is known from the nonrelativistic MW implementation in the MRChem[Wind_2023] package (written in C++), increasing the number of electrons will not pose additional theoretical challenges, but only increase the required computational resources (memory in particular) as tested by our group and extensively discussed in Ref. Wind_2023.
To assess quality of the results we have performed an array of calculations by varying the following parameters in our relativistic code:
  • The integral formulation to optimize the eigenfunction switching between using the $\hat{\mathfrak{D}}$ operator and the $\hat{\mathfrak{D}}^{2}$ operator.
  • The requested precision $\delta$ from $\delta=10^{-4}$ (MW4) to $\delta=10^{-8}$ (MW8).
  • The derivative operator: ABGV and BS
  • The placement of the nucleus: centered in the box, to be exactly on a nodal point or slightly off-centered to move it away from a nodal point.
The length of the box $L$ for the MRA was dependent on the nuclear charge: $L=\lceil\frac{50}{Z}\rceil$, to make sure the electronic density is practically zero at the boundary. The threshold for SCF convergence was set to $\epsilon=10\,\delta$. The order of the polynomial was set to $k=3-\log_{10}(\delta)$.
Convergence is achieved once the norm of the orbital update drops below $\epsilon$.
4.1 One-electron calculations
For each of the selected nuclei, we have optimized the wave function and computed the final error in the energy compared to the reference provided by GRASP[grasp_ref]. The results have been collected in four different graphs.
Figure 1: Relative error in the total energy of the H atom computed with MWs, with respect to the GRASP[grasp_ref] reference value. Each graph represents a choice of derivative operator (ABGV or BS) and atom placement (at a dyadic point or not). Each line in the graph represents a given choice of algorithm ($\Phi_{\hat{\mathfrak{D}}}$ or $\Phi_{\hat{\mathfrak{D}^{2}}}$) and expectation value ($\hat{\mathfrak{D}}$ or $\hat{\mathfrak{D}^{2}}$). Each point is the requested precision (from MW4 to MW8) of the calculation.
[…]
Each graph corresponds to one specific choice of derivative operator (ABGV or BS) and nucleus collocation (centered or off-centered with respect to the dyadic points of the grid). Each line in the graph corresponds to a specific pair of (a) optimization algorithm ($\Phi_{\hat{\mathfrak{D}}}$ in light blue and $\Phi_{\hat{\mathfrak{D}}^{2}}$ in dark blue) and (b) computation of the expectation value ($\hat{\mathfrak{D}}$ and $\hat{\mathfrak{D}^{2}}$). Solid lines correspond to diagonal setups in Eq. 31, whereas dashed lines are the off-diagonal ones. For each line in the graph, different points correspond to a different requested precision from MW4 ($\delta=10^{-4}$) to MW8 ($\delta=10^{-8}$).
@@ 4.1 One-electron calculations
Let us define the error of a given calculation as $\Delta\varepsilon(\hat{\mathcal{O}}_{E}=|\varepsilon(\hat{\mathcal{O}}_{E},\Phi_{\hat{\mathcal{O}}_{SCF}})-\varepsilon_{ref}|$, using the corresponding GRASP[grasp_ref] result as a reference.
For what concerns the diagonal entries of (31) and their respective errors, we observe that:
@@
Since no clear pattern emerged for the off-diagonal terms, we’ll only comment the results regarding the diagonal ones.
The choice of the derivative (ABGV vs BS) does not seem to affect the $\hat{\mathfrak{D}}^{2}$ calculations significantly. This is consistent with the fact that the only place where a derivative operator is applied in the $\hat{\mathfrak{D}}^{2}$ approach is in the anticommutator term in Equation 21, prior to the convolution with the Helmholtz propagator. For the algorithm using the $\hat{\mathfrak{D}}$ operator, the derivative is instead applied after the convolution, hence the B-spline version of the derivative leads to smoother functions in turns increasing precision as can be seen from Fig. 1 and Fig. 2, in which the relative error (with respect to the GRASP[grasp_ref] reference) attained at MW8 is smaller compared to the same conditions using the ABGV derivative. However, for tight thresholds (MW7-MW8), the choice of derivative has limited effect on the final precision for the $\hat{\mathfrak{D}}^{2}$ algorithm. Putting the nucleus in a dyadic or a non-dyadic point, showed no clear effect. For this reason the results of charging this particular variable are inconclusive.
In summary, the $\hat{\mathfrak{D}}^{2}$ based algorithm generally showed an improvement in the precision obtained, with gains between 1 and 4 orders of magnitude. In a few cases, the results attained by using the two algorithms yielded roughly the same precision.
4.2 Two-electron calculations
@@
We employed the instantaneous Dirac-Coulomb Hamiltonian, neglecting retardation effects of the Breit Hamiltonian and we exploited Kramers’ Time Reversal Symmetry (KTRS). With these limitations, adding a second electron does not require the computation of a second spinor explicitly, and the electron-electron interaction requires only the inclusion of the direct term $J$ for a single electron and no exchange $K$: by virtue of KTRS, the exchange term is in this case identical to the direct term, in analogy with a closed-shell non relativistic system.
Figure 3: Relative error in the total energy of the He atom computed with MWs, with respect to the GRASP[grasp_ref] reference value. Each graph represents a choice of derivative operator (ABGV or BS) and atom placement (at a dyadic point or not). Each line in the graph represents a given choice of algorithm ($\Phi_{\hat{\mathfrak{D}}}$ or $\Phi_{\hat{\mathfrak{D}^{2}}}$) and expectation value ($\hat{\mathfrak{D}}$ or $\hat{\mathfrak{D}^{2}}$). Each point is the requested precision (from MW4 to MW8) of the calculation.
@@
As for the one-electron case, the ($\Phi_{\hat{\mathfrak{D}}^{2}}$, $\hat{\mathfrak{D}}^{2}$) combination is generally more precise than the corresponding ($\Phi_{\hat{\mathfrak{D}}}$, $\hat{\mathfrak{D}}$) one, with the exception of Mercury as shown in figure 4. Moreover, the precision gain is far greater for lighter atoms, where the ($\Phi_{\hat{\mathfrak{D}}^{2}}$, $\hat{\mathfrak{D}}^{2}$) results are up to four orders of magnitude more precise than the ($\Phi_{\hat{\mathfrak{D}}}$, $\hat{\mathfrak{D}}$) ones. For heavier nuclei, such as Hg, the precision gain observed by converging the wave function with $\hat{\mathfrak{D}}^{2}$ was generally less impressive, in the range of a single order of magnitude, or less.
For two-electron systems and heavier nuclei, it has also been necessary to apply a dampening of each iteration to achieve convergence: instead of using the result of the propagator equations (15) and (22), at each iteration we have taken a liner combination of the wave functions before and after the propagation.
To understand the observed behavior for two-electron systems with heavier nuclei, we performed a series of calculations on Ar with an artificial light speed $c^{\prime}=30.83309764$ as described in Sec. 3.3 These calculations required the same dampening strategy as for Hg but provided the same precision obtained with Ar and the standard speed of light. We conclude therefore that the deeper refinement needed do describe heavier nuclei is the limiting factor in the precision attained, rather than the choice of method ($\hat{\mathfrak{D}}$ vs $\hat{\mathfrak{D}}2$). On the other hand, the magnitude of the relativistic correction affects the convergence of the calculation.
@@
To investigate this further, we attempted to run some calculations using a Fermi-Dirac nuclear distribution for an artificial Hg nucleus with a radius 10 and 100 times the regular one to achieve a smoother potential, requiring a less aggressive refinement in the vicinity of the nucleus. However, none of these calculations converged.
4.3 Performance considerations
Although the $\hat{\mathfrak{D}}^{2}$-based algorithm is promising to achieve precise results, it requires a larger amount of memory, and each iteration is slower than the corresponding $\hat{\mathfrak{D}}-$based algorithm. This can be traced back to the need to compute more terms in the potential. Moreover, computing functions like $\hat{V}^{2}\Phi_{i}$, is numerically demanding because it requires a finer refinement of the grid close to the nucleus, further increasing the memory footprint and the time required for each iteration. This is especially evident for heavier nuclei where this is also the primary factor limiting the precision obtained. As noted above, two-electron systems also require dampening to reach convergence.
We underline that this is only a prototype implementation. The tradeoff between precision and performance will guide the design choices for a production-grade implementation. In such an implementation, it will moreover be possible to push precision requirements to achieve better results for heavier nuclei. Finally, the need for dampening will be superseded by off-the-shelf accelerators such as Direct Inversion in the Iterative Subspace (DIIS[Pulay1980-xh] or Krylov-accelerated inexact Newton (KAIN[Harrison101002jcc10108].
A further possible performance gain for the $\hat{\mathfrak{D}}^{2}$-based algorithm, especially when working with larger molecular systems, would be to apply the different parts of the generalized potential in Eq. 21 with a threshold scaled by their respective order in $c$.
5 Conclusions
We have developed and implemented the squared Dirac operator, $\hat{\mathfrak{D}}^{2}$, in a MW framweork, following the original formulation by Kutzelnigg and Wallmeier[Kutzelnigg_1984, Wallmeier_1981]. The main motivation was to obtain a variationally well-posed relativistic mean-field scheme. By squaring the Dirac operator, the negative-energy continuum is folded onto the positive-energy branch, resulting in an operator that is bounded from below. This transformation enables genuine energy minimization and systematic numerical convergence. However, realizing these advantages in practice requires a numerical representation that does not introduce uncontrolled basis-set errors. A MW basis is crucial for this strategy. Unlike conventional atom-centered basis sets, MWs provide a near-complete representation within a prescribed tolerance, which is essential when matrix elements of $\hat{\mathfrak{D}}^{2}$ are expressed as products (see Eq. (23))
Furthermore, the squared formulation leads to an integral equation that closely resembles the non-relativistic Helmholtz-propagator strategy, allowing established multiresolution algorithms to be reused with limited modifications while retaining a four-component description.
@@ 5 Conclusions
In this implementation, $\hat{\mathfrak{D}}^{2}$ is less sensitive to the derivative representation (ABGV vs. B-spline), and no systematic dependence on nuclear placement relative to the grid is observed, especially for tight thresholds (MW8). For two-electron systems and heavier nuclei, stable convergence required damped orbital updates.
The key drawback of our implementation is only computational, because the generalized potential $\hat{\mathcal{V}}$ introduces extra terms with notably repeated application of $V$ (effectively $V^{2}\Phi$), which increases near-nucleus refinement, memory, and wall time. This was seen specifically for heavy nuclei where refinement already limits attainable precision.
Future work will address how to to migrate from developing version to a production-grade parallel implementation with tighter thresholding, add robust SCF acceleration to remove ad hoc dampening, it will also extend the formulation to general many-electron molecules and relativistic Hamiltonians beyond the Dirac–Coulomb picture.
Data Availability
All data supporting the findings of this study are provided within the article and the Supporting Information. The code used to generate the results and figures is publicly available at the ReMRChem[ReMRChem] github repository.